Fermi National Laboratory
 

Accelerator Update for August 3, 2007 - August 6, 2007

 

Friday August 3

The day shift began with Operations monitoring the Tevatron (TeV) store 5590 (11.1E30) and the Recycler stash (217.3E10), with the Antiproton Source (Pbar) stacking (171.1E10), with Linac, Booster, and Main Injector (MI) operating normally, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams, and with MiniBooNE taking beam. NuMI will remain off until after the summer shutdown.

Operations began preparing Booster and Pbar Pre-Vault area for access at 9:04 AM.

At 10:38 AM, a Linac RF station (LRF3) tripped off. It reset. Also, LRF4 tripped off for the first of four times in about an hour. A Linac expert investigated after the last trip and brought the station back on line.

At 10:50 AM, LRF3 tripped off. It reset.

Booster came back on line at 12:39 PM.

Operations resumed sending beam to MI and MiniBooNE at 12:41 PM.

Operations established beam to the Meson MTest experiment at 12:58 PM.

At 1:04 PM, Booster suffered vacuum bursts that tripped off vacuum pumps and a power supply (GMPS). The vacuum pumps were reset.

At 1:17 PM, BRF17 tripped off. It reset.

MiniBooNE reported at 1:19 PM that they had turned off their horn.

At 1:36 PM, Operations turned off Booster and prepared it for an access to investigate vacuum problems.

At 2:29 PM. LRF5 tripped off. It reset.

Pelletron experts accessed the MI-31 area at 2:41 PM. They were out of access about 75 minutes later.

At 3:19 PM, Operations received a report of water bubbling up out of the ground near the Booster East Anode Power Supply. Operators contacted experts and had the area roped off.

Operations reported at 4:32 PM that the Booster vacuum leak went away before it was found.

Booster resumed sending beam at 6:09 PM.

At 6:18 PM, Operations reported that due to a vacuum leak in MP02 MiniBooNE/SciBooNE was offline and would not be back up until after the summer shutdown ends. (Later, operators discovered the leak was caused by the Booster High Rep rate studies and went away when the studies ended.)

At 6:19 PM, Pelletron experts accessed MI-31 to investigate a problem.

Operations prepared MiniBooNE for access at 8:57 PM.

At 11:25 PM, LRF4 tripped off several times over the next 20 minutes. It reset.

Operations reported at 11:52 PM that the Recycler had stashed antiprotons once during the evening shift.

Saturday August 4

The midnight shift began with Operations monitoring store 5590 (30.15E30) and the stash (230.2E10), with Pbar stacking (19.6E10), and with MiniBooNE and MTest taking beam.

Operations disabled the MTest beam at 12:01 AM.

At 4:27 AM, BRF11 tripped off for the first of two times in 12 minutes. It reset.

Operations terminated store 5590 at 5:06 AM.

TeV experts began a study period at 5:18 AM.

At 6:06 AM, Operations reported problems with all of the Linac quadrupole power supplies.

MTest resumed taking beam at 6:12 AM.

Operations reported at 6:28 AM that the Linac quadrupoles were tripping off due to temperature problems. The Duty Mechanic investigated and reported the air conditioning in the Preaccelerator control room was running full out.

Per a Linac expert’s suggestion, operators lowered the current for IQTM1U by 8 amps; operators reported that this seemed to resolve the quadrupole problems.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

The TeV studies ended at 8:40 AM.

Operations put the TeV through a dry squeeze at 8:46 AM.

Operations put the TeV into shot setup at 9:15 AM.

At 9:30 AM, a Booster vacuum pump tripped off (IPRF17) and caused BRF7 & 8 to trip off. Everything reset.

At 10:58 AM, a Linac Klystron RF station (KRF5) tripped off. It reset.

At 11:15 AM, Operations reported that the third transfer of antiprotons to the TeV didn’t make it. Operators investigated and discovered a corrector magnet (VT901) was not set to its correct value. This was corrected and all the other transfers ended up making it okay.

Operations established store 5592 at 11:56 AM with an initial luminosity of 215.64E30.

A Pbar expert began a one-hour study at 12:06 PM.

Operators began preparing Pbar for the summer shutdown at 2:28 PM.

A Recycler expert began a 15-minute MI no-ramp study at 2:39 PM.

Operations disabled the MTest beam at 6:01 PM.

At 6:43 PM, the Duty Electrician reported that the Pbar AP50 emergency generator was running and had been for about four hours. He investigated and discovered that something was off that required the generator to be on, but he couldn’t discover what it was. Operations investigated and then called in an expert.

A Recycler expert began another MI no-ramp study at 7:29 PM.

Operators locked off Switchyard at 7:29 PM in preparation for the summer shutdown.

Operations terminated store 5592 at 8:03 PM.

Recycler studies ended at 8:14 PM and MI resumed ramping.

TeV experts began a study period at 8:21 PM.

At 8:38 PM, LRF3 tripped off for the first of two times in five minutes. It reset.

At 8:48 PM, the TeV suffered a quench at sector D3, and other areas.

At 9:29 PM, LRF4 tripped off. It reset.

The cryo system technicians at CHL volunteered at 9:44 PM to conduct the TeV quench recovery.

Sunday August 5

The midnight shift began with TeV studies on hold while the TeV recovered from a quench.

Operations put the TeV through a dry squeeze at 12:46 AM.

TeV studies resumed at 1:02 AM.

Operations began preparing Booster, MI-8, and MI for access at 1:08 AM.

Booster and MI came out of access and at 2:30 AM, Operations prepared the Booster for a ‘high repetition rate of beam’ study.

At 3:16 AM, BRF9 tripped off. It reset.

TeV studies ended at 3:27 AM.

Operations prepared MI for access at 3:51 AM.

Operators began conducting a hipot of the TeV at 3:52 AM.

Operations reported at 4:49 AM that they had switched off feeder 49.

Operations reported at 4:51 AM that all the sector vacuum valves in the TeV and MI had been closed.

At 5:55 AM, LRF4 tripped off. It reset.

By 8 AM, Operations had prepared the TeV, MI, the MI-8 transfer line, and Booster for accesses.

Cryo system experts reported at 8:25 AM that they were beginning to warm up the TeV cryo systems at C4 and D1. The warm up took about six and a half hours.

Booster RF experts began work on the RF system at 9:34 AM.

Operations reported at noon that a scheduled four-hour Kautz Road Substation (KRSS) power outage had begun. The MI and Recycler areas were being back fed from the Master Substation (MSS) during this outage.

A Linac expert began a study period at 3:14 PM.

At 5:10 PM, Operations reported that the ComEd KRSS work had been delayed due to rain.

Operations reported that the ComEd KRSS work had been completed. The power outage to switch MI back to the Kautz Road Substation began at 8:53 PM and lasted about 10 minutes.

The Linac studies ended at 10:28 PM.

Operations reported at 11:59 PM that there were problems recovering the LCW and vacuum for MI, NuMI, MiniBooNE, the Pelletron, and the Recycler after the last power outage.

Monday August 6

The midnight shift began with Operations trying to recover systems from the scheduled power outage. Operations called in experts to help with the recoveries.

Operations reported that a scheduled 30-minute Master substation power outage began at 7 AM.

Operations reported at 7:58 AM that operators and experts had worked at recovering systems during most of shift.

The Plans for Monday

The plans for today and for the next 10 weeks are to begin all planned shutdown work. Neutron Therapy Facility will run during the evenings through out the shutdown except for one 10-day stretch. This is the last of the normal Accelerator Updates for the duration of the shutdown. A Shutdown Update will be posted once a week during the shutdown.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/06/2007    email Fermilab